# When Pre-trained Visual Representations Fall Short: Limitations in Visuo-Motor Robot Learning

### v1

@@ Appendix B On the use of PVRs in Downstream Robotics Tasks
For instance, in manipulation tasks utilizing dense visual descriptors without a learnable policy component, the field has transitioned from training these features from scratch (Sundaresan et al., 2020; Ganapathi et al., 2021; Manuelli et al., 2021; Florence et al., 2019, 2018; Yen-Chen et al., 2022) to leveraging features extracted from vision foundation models such as DINO (Caron et al., 2021; Oquab et al., 2024) and CLIP (Radford et al., 2021). A common strategy involves integrating these features into 3D spatial representations, as exemplified by works like F3RM (Shen et al., 2023), D3Fields (Wang et al., 2024), and Click2Grasp (Tsagkas et al., 2024). Similarly, PVRs have significantly contributed to semantic mapping, particularly when incorporating language components, as demonstrated in VL-Fields (Tsagkas et al., 2023) and CLIP-Fields (Shafiullah et al., 2023).
However, PVRs have yet to gain traction in robot learning frameworks. The foundational study by (Chi et al., 2023) shows that diffusion policies conditioned on R3M (Nair et al., 2022) features perform worse than those conditioned on visual encoders trained end-to-end. Notably, policies trained with PVRs tend to produce jittery actions and are prone to getting stuck, which may be linked to an incorrect assumption of the Markov property. Consequently, end-to-end training of visual encoders alongside the policy remains the preferred approach (e.g., (Sochopoulos et al., 2024; Florence et al., 2021)).
Finally, an interesting side that is out of the scope of this research concerns the use of other modalities in robot learning. More specifically, a group of recent works has shifted from PVRs that utilise only image data but also language (Li et al., 2023; Gupta et al., 2024; Ma et al., 2023a; Karamcheti et al., 2023).
@@ Appendix C Temporal Encoding
Figs. A1 and A2 extend the methodology from Fig. 1 to real-world data from (Chi et al., 2023; Zhou et al., 2023a, b). As shown, similar issues arise, such as the temporal entanglement of features. While testing our approach on a real robot remains part of our future work, these results provide evidence that the challenges identified in simulation persist in real-world scenarios. Although it is well established that simulations like MetaWorld serve as a reliable proxy for real-world performance (Silwal et al., 2024), these figures underscore the significance of our findings and the necessity of our proposed solutions.
@@ C.1 On the quality of simulation data
Figure A1: Visualisation of PCA results for the features extracted from frames of an expert demonstration trajectory from the real-world in the Push-T task (Chi et al., 2023) across the studied PVRs. In this example, we see mostly a cluster forming when the robot has pushed the T block to its outline and makes very small adjustments.
Figure A2: Visualisation of PCA results for the features extracted from frames of an expert demonstration trajectory from the real-world in the ASU Table (Zhou et al., 2023a, b) task across the studied PVRs. In this example, entanglement occurs mostly from the fact that the robot picks up the can, rotates it by 90 degrees and then places it in the exact same spot.
@@ C.2 TE Per Task Results
T DINOv2 DINOv1 MAE CLIP ViT iBot VC1 MoCov2 SWAV VIP DenseCL R3M VFS VICRegL Bin Picking -- $43.3\pm 0.5$ $33.3\pm 2.4$ $50.0\pm 4.3$ $56.3\pm 1.9$ $26.7\pm 5.2$ $41.7\pm 3.3$ $52.0\pm 2.9$ $56.7\pm 7.4$ $54.3\pm 3.3$ $32.7\pm 3.3$ $40.7\pm 3.4$ $40.3\pm 5.4$ $51.0\pm 4.1$ $46.0\pm 7.8$triangledown $28.3\pm 3.1$ $35.3\pm 2.4$ $18.7\pm 2.1$ $46.7\pm 4.8$ $32.7\pm 5.2$ $22.0\pm 2.2$ $45.0\pm 4.9$ $38.3\pm 9.6$ $42.3\pm 0.5$ $37.7\pm 7.1$ $29.7\pm 1.7$ $49.7\pm 6.5$ $62.3\pm 3.1$ $44.7\pm 10.4$diamond $66.3\pm 5.2$ $53.7\pm 3.4$ $53.7\pm 0.9$ $71.0\pm 2.2$ $56.0\pm 2.4$ $53.3\pm 4.9$ $53.7\pm 9.9$ $65.7\pm 3.8$ $63.7\pm 4.6$ $51.7\pm 2.1$ $40.0\pm 8.0$ $77.0\pm 6.7$ $79.7\pm 0.9$ $68.7\pm 0.5$ Disassemble -- $85.7\pm 0.5$ $77.3\pm 0.5$ $35.7\pm 1.9$ $76.7\pm 4.2$ $81.0\pm 2.2$ $90.3\pm 3.1$ $74.7\pm 2.9$ $92.3\pm 1.2$ $91.3\pm 0.5$ $85.3\pm 2.5$ $87.3\pm 2.9$ $83.0\pm 3.6$ $80.0\pm 1.6$ $89.0\pm 0.8$triangledown $91.7\pm 1.7$ $88.7\pm 4.0$ $76.7\pm 2.5$ $72.7\pm 3.4$ $82.0\pm 1.4$ $89.0\pm 2.4$ $84.7\pm 3.7$ $91.3\pm 0.5$ $91.0\pm 2.2$ $24.0\pm 7.3$ $87.3\pm 1.7$ $91.3\pm 3.9$ $92.0\pm 0.8$ $89.7\pm 2.5$diamond $93.7\pm 0.9$ $91.0\pm 1.6$ $92.0\pm 2.2$ $83.3\pm 2.9$ $90.7\pm 1.7$ $95.3\pm 2.1$ $77.3\pm 0.5$ $92.0\pm 0.8$ $94.3\pm 0.9$ $91.0\pm 2.2$ $90.0\pm 0.8$ $91.0\pm 2.9$ $93.7\pm 2.5$ $93.0\pm 1.6$ Coffee Pull -- $46.0\pm 3.6$ $53.3\pm 2.1$ $42.3\pm 3.3$ $42.0\pm 0.8$ $49.3\pm 1.2$ $55.7\pm 0.5$ $19.3\pm 7.1$ $72.3\pm 1.9$ $72.0\pm 0.8$ $99.7\pm 0.5$ $62.3\pm 1.7$ $89.0\pm 2.4$ $89.3\pm 3.8$ $82.7\pm 5.0$triangledown $47.3\pm 5.4$ $50.7\pm 1.7$ $63.0\pm 5.0$ $50.3\pm 2.1$ $49.0\pm 1.6$ $57.0\pm 1.6$ $60.3\pm 4.2$ $77.0\pm 0.8$ $77.0\pm 2.2$ $88.3\pm 1.7$ $59.3\pm 0.5$ $99.0\pm 0.8$ $86.3\pm 3.3$ $74.0\pm 2.2$diamond $56.0\pm 3.3$ $54.0\pm 2.4$ $47.7\pm 4.7$ $51.3\pm 3.3$ $45.0\pm 1.4$ $58.0\pm 2.9$ $61.0\pm 3.7$ $72.0\pm 2.4$ $77.0\pm 3.6$ $99.0\pm 0.8$ $63.0\pm 3.7$ $93.7\pm 2.6$ $92.7\pm 2.6$ $79.3\pm 2.5$ Shelf Place -- $9.0\pm 4.2$ $6.0\pm 2.2$ $5.0\pm 0.8$ $10.3\pm 0.9$ $1.3\pm 1.9$ $4.0\pm 0.0$ $2.7\pm 1.7$ $17.0\pm 4.5$ $28.7\pm 1.2$ $28.7\pm 4.5$ $8.0\pm 0.8$ $39.7\pm 4.8$ $20.3\pm 3.3$ $20.7\pm 4.2$triangledown $8.0\pm 0.8$ $6.7\pm 3.1$ $8.7\pm 1.2$ $9.3\pm 0.5$ $4.3\pm 1.9$ $3.7\pm 1.9$ $1.7\pm 0.5$ $22.7\pm 1.7$ $23.3\pm 3.4$ $32.0\pm 1.6$ $15.0\pm 4.2$ $43.7\pm 3.4$ $21.0\pm 1.6$ $27.7\pm 2.6$diamond $3.7\pm 2.1$ $5.0\pm 1.4$ $4.7\pm 2.9$ $12.3\pm 1.7$ $6.7\pm 2.5$ $4.0\pm 0.8$ $4.3\pm 2.1$ $20.7\pm 4.2$ $33.7\pm 5.2$ $22.0\pm 2.9$ $12.3\pm 2.1$ $33.3\pm 1.2$ $26.7\pm 2.9$ $28.0\pm 0.8$ Peg Insert Side -- $30.7\pm 1.2$ $23.7\pm 2.1$ $23.0\pm 1.4$ $18.7\pm 3.7$ $4.3\pm 1.2$ $20.7\pm 3.4$ $22.0\pm 1.4$ $34.0\pm 0.8$ $48.3\pm 3.4$ $34.0\pm 2.2$ $32.3\pm 3.7$ $59.7\pm 0.5$ $38.0\pm 5.7$ $38.7\pm 2.1$triangledown $24.3\pm 3.8$ $33.0\pm 5.4$ $28.3\pm 4.7$ $13.7\pm 4.9$ $12.0\pm 2.8$ $25.3\pm 1.7$ $17.0\pm 2.4$ $35.7\pm 3.7$ $39.3\pm 6.6$ $44.7\pm 6.3$ $23.3\pm 2.4$ $34.7\pm 5.2$ $35.7\pm 0.9$ $43.7\pm 3.4$diamond $30.7\pm 0.9$ $31.3\pm 2.4$ $27.3\pm 2.9$ $18.3\pm 0.5$ $25.7\pm 1.7$ $11.0\pm 2.8$ $8.7\pm 0.5$ $50.7\pm 0.5$ $37.3\pm 1.7$ $74.0\pm 0.8$ $40.3\pm 3.3$ $69.0\pm 0.8$ $50.3\pm 3.4$ $44.3\pm 3.4$ Box Close -- $44.0\pm 2.2$ $59.3\pm 5.4$ $63.0\pm 6.7$ $46.3\pm 5.4$ $57.7\pm 7.4$ $61.7\pm 1.7$ $56.3\pm 6.1$ $56.3\pm 5.3$ $53.0\pm 2.2$ $48.7\pm 6.2$ $59.7\pm 4.5$ $49.0\pm 6.2$ $57.3\pm 1.2$ $61.3\pm 4.5$triangledown $45.0\pm 0.8$ $61.0\pm 1.6$ $64.0\pm 7.8$ $55.0\pm 1.6$ $54.0\pm 2.2$ $64.7\pm 5.6$ $60.3\pm 2.1$ $58.7\pm 5.4$ $57.7\pm 4.8$ $33.3\pm 18.8$ $57.3\pm 5.3$ $41.3\pm 2.6$ $62.0\pm 1.6$ $67.3\pm 2.5$diamond $56.7\pm 1.2$ $68.3\pm 0.5$ $57.0\pm 2.2$ $63.3\pm 5.0$ $73.3\pm 1.7$ $61.0\pm 5.0$ $71.0\pm 2.2$ $67.0\pm 5.0$ $61.0\pm 0.8$ $52.7\pm 0.5$ $62.3\pm 3.7$ $68.0\pm 2.2$ $72.0\pm 3.7$ $68.3\pm 3.3$ Assembly -- $67.3\pm 3.7$ $61.7\pm 2.9$ $7.3\pm 0.9$ $35.3\pm 2.5$ $55.7\pm 2.5$ $72.7\pm 3.8$ $82.3\pm 0.5$ $90.3\pm 3.1$ $94.3\pm 0.5$ $5.0\pm 3.6$ $97.0\pm 2.2$ $93.3\pm 1.2$ $92.3\pm 1.7$ $93.0\pm 2.4$triangledown $24.7\pm 0.5$ $47.3\pm 2.9$ $59.7\pm 2.1$ $34.3\pm 2.5$ $39.7\pm 3.8$ $63.3\pm 0.9$ $50.3\pm 2.1$ $82.3\pm 4.7$ $93.0\pm 2.2$ $35.7\pm 29.8$ $77.0\pm 2.9$ $98.3\pm 2.4$ $92.7\pm 2.1$ $93.3\pm 1.2$diamond $73.0\pm 4.1$ $79.3\pm 3.1$ $66.0\pm 4.3$ $53.7\pm 1.9$ $79.3\pm 2.1$ $61.3\pm 2.5$ $63.3\pm 3.1$ $90.3\pm 4.5$ $91.0\pm 2.8$ $71.7\pm 8.7$ $93.7\pm 3.4$ $100.0\pm 0.0$ $87.7\pm 3.3$ $92.7\pm 2.9$ Button Press Wall -- $70.3\pm 4.5$ $76.0\pm 2.9$ $49.3\pm 1.9$ $76.7\pm 4.2$ $62.3\pm 9.0$ $57.7\pm 1.2$ $31.0\pm 0.0$ $57.3\pm 0.9$ $69.3\pm 2.6$ $22.3\pm 17.4$ $52.0\pm 7.9$ $63.3\pm 3.9$ $58.7\pm 4.6$ $51.3\pm 5.4$triangledown $59.3\pm 7.0$ $84.7\pm 2.9$ $63.3\pm 4.9$ $67.7\pm 1.2$ $64.7\pm 1.2$ $74.0\pm 5.4$ $65.0\pm 0.8$ $63.7\pm 7.5$ $64.0\pm 4.9$ $65.7\pm 14.8$ $68.0\pm 2.9$ $60.3\pm 2.6$ $62.0\pm 6.5$ $50.0\pm 1.6$diamond $79.0\pm 4.2$ $77.3\pm 2.4$ $85.0\pm 0.0$ $82.0\pm 2.8$ $69.0\pm 4.3$ $74.3\pm 0.9$ $63.0\pm 3.7$ $77.7\pm 8.7$ $64.7\pm 3.4$ $6.3\pm 2.5$ $83.3\pm 4.8$ $54.7\pm 1.2$ $58.0\pm 5.7$ $67.0\pm 12.7$ Pick Place Wall -- $39.7\pm 7.4$ $29.3\pm 4.5$ $41.0\pm 4.5$ $36.0\pm 5.7$ $59.3\pm 7.3$ $25.7\pm 1.7$ $49.7\pm 7.1$ $84.0\pm 1.6$ $69.3\pm 1.7$ $54.0\pm 15.7$ $61.7\pm 5.4$ $67.0\pm 4.9$ $80.7\pm 0.5$ $76.3\pm 2.4$triangledown $47.0\pm 6.2$ $69.3\pm 5.3$ $52.0\pm 4.2$ $49.0\pm 2.2$ $64.3\pm 3.3$ $58.7\pm 2.1$ $44.3\pm 2.6$ $88.7\pm 2.4$ $89.0\pm 0.8$ $86.3\pm 2.5$ $59.0\pm 9.6$ $95.3\pm 0.5$ $90.0\pm 0.8$ $95.0\pm 2.4$diamond $24.7\pm 2.5$ $59.0\pm 7.3$ $26.7\pm 4.5$ $29.7\pm 4.2$ $40.0\pm 2.2$ $24.3\pm 0.9$ $26.0\pm 4.3$ $79.7\pm 0.9$ $62.0\pm 3.7$ $86.7\pm 2.6$ $49.7\pm 2.6$ $66.0\pm 0.8$ $83.3\pm 5.2$ $73.3\pm 4.6$ Drawer Open -- $90.7\pm 1.2$ $92.0\pm 0.0$ $92.7\pm 1.2$ $87.7\pm 0.9$ $89.3\pm 0.5$ $96.3\pm 1.2$ $95.7\pm 1.7$ $93.3\pm 1.9$ $93.3\pm 1.9$ $95.7\pm 1.7$ $100.0\pm 0.0$ $90.3\pm 1.2$ $95.0\pm 0.8$ $92.0\pm 0.0$triangledown $93.0\pm 1.4$ $93.3\pm 1.9$ $94.0\pm 1.6$ $90.3\pm 1.2$ $89.3\pm 0.5$ $99.7\pm 0.5$ $100.0\pm 0.0$ $95.0\pm 0.8$ $91.7\pm 0.5$ $99.3\pm 0.9$ $100.0\pm 0.0$ $100.0\pm 0.0$ $95.0\pm 0.8$ $92.0\pm 0.0$diamond $100.0\pm 0.0$ $100.0\pm 0.0$ $100.0\pm 0.0$ $99.3\pm 0.9$ $99.3\pm 0.9$ $100.0\pm 0.0$ $100.0\pm 0.0$ $95.7\pm 0.5$ $100.0\pm 0.0$ $95.7\pm 1.7$ $98.0\pm 0.8$ $100.0\pm 0.0$ $100.0\pm 0.0$ $94.7\pm 0.9$

### v2

@@ Appendix A Evaluated PVRs
A.1 The Importance of Training Datasets
The dataset(s) used for pre-training plays a pivotal role in PVRs. While it was hypothesised that pre-training with video data featuring egocentric human-object interaction would be highly effective for learning features suitable for robot learning (due to their emphasis on object manipulation), research indicates that the diversity of images within the dataset is a more critical factor in successful robot learning [54, 8]. Indeed, PVRs pre-trained on static datasets such as ImageNet [55] have demonstrated competitive performance, underscoring the importance of dataset variability over modality.
@@ Appendix B On the use of PVRs in Downstream Robotics Tasks
For instance, in manipulation tasks utilizing dense visual descriptors without a learnable policy component, the field has transitioned from training these features from scratch [56, 57, 58, 59, 60, 61] to leveraging features extracted from vision foundation models such as DINO [10, 44] and CLIP [45]. A common strategy involves integrating these features into 3D spatial representations, as exemplified by works like F3RM [5], D3Fields [62], and Click2Grasp [4]. Similarly, PVRs have significantly contributed to semantic mapping, particularly when incorporating language components, as demonstrated in VL-Fields [63] and CLIP-Fields [64].
However, PVRs have yet to gain traction in robot learning frameworks. The foundational study by [13] shows that diffusion policies conditioned on R3M [27] features perform worse than those conditioned on visual encoders trained end-to-end. Notably, policies trained with PVRs tend to produce jittery actions and are prone to getting stuck, which may be linked to an incorrect assumption of the Markov property. Consequently, end-to-end training of visual encoders alongside the policy remains the preferred approach (e.g., [65, 66]).
Finally, an interesting side that is out of the scope of this research concerns the use of other modalities in robot learning. More specifically, a group of recent works has shifted from PVRs that utilise only image data but also language [67, 68, 69, 70].
@@ Appendix C Temporal Encoding
Figs. A1 and A2 extend the methodology from Fig. 1 to real-world data from [13, 71, 72]. As shown, similar issues arise, such as the temporal entanglement of features. While testing our approach on a real robot remains part of our future work, these results provide evidence that the challenges identified in simulation persist in real-world scenarios. Although it is well established that simulations like MetaWorld serve as a reliable proxy for real-world performance [18], these figures underscore the significance of our findings and the necessity of our proposed solutions.
@@ C.1 On the Quality of Simulation Data
Figure A1: Visualisation of PCA results for the features extracted from frames of an expert demonstration trajectory from the real-world in the Push-T task [13] across the studied PVRs. In this example, we see mostly a cluster forming when the robot has pushed the T block to its outline and makes very small adjustments.
Figure A2: Visualisation of PCA results for the features extracted from frames of an expert demonstration trajectory from the real-world in the ASU Table [71, 72] task across the studied PVRs. In this example, entanglement occurs mostly from the fact that the robot picks up the can, rotates it by 90 degrees and then places it in the exact same spot.
@@ C.2 TE Per Task Results
T DINOv2 DINOv1 MAE CLIP ViT iBot VC1 MoCov2 SWAV VIP DenseCL R3M VFS VICRegL Bin Picking -- $43.3\pm 0.5$ $33.3\pm 2.4$ $50.0\pm 4.3$ $56.3\pm 1.9$ $26.7\pm 5.2$ $41.7\pm 3.3$ $52.0\pm 2.9$ $56.7\pm 7.4$ $54.3\pm 3.3$ $32.7\pm 3.3$ $40.7\pm 3.4$ $40.3\pm 5.4$ $51.0\pm 4.1$ $46.0\pm 7.8$triangledown $28.3\pm 3.1$ $35.3\pm 2.4$ $18.7\pm 2.1$ $46.7\pm 4.8$ $32.7\pm 5.2$ $22.0\pm 2.2$ $45.0\pm 4.9$ $38.3\pm 9.6$ $42.3\pm 0.5$ $37.7\pm 7.1$ $29.7\pm 1.7$ $49.7\pm 6.5$ $62.3\pm 3.1$ $44.7\pm 10.4$diamond $66.3\pm 5.2$ $53.7\pm 3.4$ $53.7\pm 0.9$ $71.0\pm 2.2$ $56.0\pm 2.4$ $53.3\pm 4.9$ $53.7\pm 9.9$ $65.7\pm 3.8$ $63.7\pm 4.6$ $51.7\pm 2.1$ $40.0\pm 8.0$ $77.0\pm 6.7$ $79.7\pm 0.9$ $68.7\pm 0.5$ Disassemble -- $85.7\pm 0.5$ $77.3\pm 0.5$ $35.7\pm 1.9$ $76.7\pm 4.2$ $81.0\pm 2.2$ $90.3\pm 3.1$ $74.7\pm 2.9$ $92.3\pm 1.2$ $91.3\pm 0.5$ $85.3\pm 2.5$ $87.3\pm 2.9$ $83.0\pm 3.6$ $80.0\pm 1.6$ $89.0\pm 0.8$triangledown $91.7\pm 1.7$ $88.7\pm 4.0$ $76.7\pm 2.5$ $72.7\pm 3.4$ $82.0\pm 1.4$ $89.0\pm 2.4$ $84.7\pm 3.7$ $91.3\pm 0.5$ $91.0\pm 2.2$ $24.0\pm 7.3$ $87.3\pm 1.7$ $91.3\pm 3.9$ $92.0\pm 0.8$ $89.7\pm 2.5$diamond $93.7\pm 0.9$ $91.0\pm 1.6$ $92.0\pm 2.2$ $83.3\pm 2.9$ $90.7\pm 1.7$ $95.3\pm 2.1$ $77.3\pm 0.5$ $92.0\pm 0.8$ $94.3\pm 0.9$ $91.0\pm 2.2$ $90.0\pm 0.8$ $91.0\pm 2.9$ $93.7\pm 2.5$ $93.0\pm 1.6$ Coffee Pull -- $46.0\pm 3.6$ $53.3\pm 2.1$ $42.3\pm 3.3$ $42.0\pm 0.8$ $49.3\pm 1.2$ $55.7\pm 0.5$ $19.3\pm 7.1$ $72.3\pm 1.9$ $72.0\pm 0.8$ $99.7\pm 0.5$ $62.3\pm 1.7$ $89.0\pm 2.4$ $89.3\pm 3.8$ $82.7\pm 5.0$triangledown $47.3\pm 5.4$ $50.7\pm 1.7$ $63.0\pm 5.0$ $50.3\pm 2.1$ $49.0\pm 1.6$ $57.0\pm 1.6$ $60.3\pm 4.2$ $77.0\pm 0.8$ $77.0\pm 2.2$ $88.3\pm 1.7$ $59.3\pm 0.5$ $99.0\pm 0.8$ $86.3\pm 3.3$ $74.0\pm 2.2$diamond $56.0\pm 3.3$ $54.0\pm 2.4$ $47.7\pm 4.7$ $51.3\pm 3.3$ $45.0\pm 1.4$ $58.0\pm 2.9$ $61.0\pm 3.7$ $72.0\pm 2.4$ $77.0\pm 3.6$ $99.0\pm 0.8$ $63.0\pm 3.7$ $93.7\pm 2.6$ $92.7\pm 2.6$ $79.3\pm 2.5$ Shelf Place -- $9.0\pm 4.2$ $6.0\pm 2.2$ $5.0\pm 0.8$ $10.3\pm 0.9$ $1.3\pm 1.9$ $4.0\pm 0.0$ $2.7\pm 1.7$ $17.0\pm 4.5$ $28.7\pm 1.2$ $28.7\pm 4.5$ $8.0\pm 0.8$ $39.7\pm 4.8$ $20.3\pm 3.3$ $20.7\pm 4.2$triangledown $8.0\pm 0.8$ $6.7\pm 3.1$ $8.7\pm 1.2$ $9.3\pm 0.5$ $4.3\pm 1.9$ $3.7\pm 1.9$ $1.7\pm 0.5$ $22.7\pm 1.7$ $23.3\pm 3.4$ $32.0\pm 1.6$ $15.0\pm 4.2$ $43.7\pm 3.4$ $21.0\pm 1.6$ $27.7\pm 2.6$diamond $3.7\pm 2.1$ $5.0\pm 1.4$ $4.7\pm 2.9$ $12.3\pm 1.7$ $6.7\pm 2.5$ $4.0\pm 0.8$ $4.3\pm 2.1$ $20.7\pm 4.2$ $33.7\pm 5.2$ $22.0\pm 2.9$ $12.3\pm 2.1$ $33.3\pm 1.2$ $26.7\pm 2.9$ $28.0\pm 0.8$ Peg Insert Side -- $30.7\pm 1.2$ $23.7\pm 2.1$ $23.0\pm 1.4$ $18.7\pm 3.7$ $4.3\pm 1.2$ $20.7\pm 3.4$ $22.0\pm 1.4$ $34.0\pm 0.8$ $48.3\pm 3.4$ $34.0\pm 2.2$ $32.3\pm 3.7$ $59.7\pm 0.5$ $38.0\pm 5.7$ $38.7\pm 2.1$triangledown $24.3\pm 3.8$ $33.0\pm 5.4$ $28.3\pm 4.7$ $13.7\pm 4.9$ $12.0\pm 2.8$ $25.3\pm 1.7$ $17.0\pm 2.4$ $35.7\pm 3.7$ $39.3\pm 6.6$ $44.7\pm 6.3$ $23.3\pm 2.4$ $34.7\pm 5.2$ $35.7\pm 0.9$ $43.7\pm 3.4$diamond $30.7\pm 0.9$ $31.3\pm 2.4$ $27.3\pm 2.9$ $18.3\pm 0.5$ $25.7\pm 1.7$ $11.0\pm 2.8$ $8.7\pm 0.5$ $50.7\pm 0.5$ $37.3\pm 1.7$ $74.0\pm 0.8$ $40.3\pm 3.3$ $69.0\pm 0.8$ $50.3\pm 3.4$ $44.3\pm 3.4$ Box Close -- $44.0\pm 2.2$ $59.3\pm 5.4$ $63.0\pm 6.7$ $46.3\pm 5.4$ $57.7\pm 7.4$ $61.7\pm 1.7$ $56.3\pm 6.1$ $56.3\pm 5.3$ $53.0\pm 2.2$ $48.7\pm 6.2$ $59.7\pm 4.5$ $49.0\pm 6.2$ $57.3\pm 1.2$ $61.3\pm 4.5$triangledown $45.0\pm 0.8$ $61.0\pm 1.6$ $64.0\pm 7.8$ $55.0\pm 1.6$ $54.0\pm 2.2$ $64.7\pm 5.6$ $60.3\pm 2.1$ $58.7\pm 5.4$ $57.7\pm 4.8$ $33.3\pm 18.8$ $57.3\pm 5.3$ $41.3\pm 2.6$ $62.0\pm 1.6$ $67.3\pm 2.5$diamond $56.7\pm 1.2$ $68.3\pm 0.5$ $57.0\pm 2.2$ $63.3\pm 5.0$ $73.3\pm 1.7$ $61.0\pm 5.0$ $71.0\pm 2.2$ $67.0\pm 5.0$ $61.0\pm 0.8$ $52.7\pm 0.5$ $62.3\pm 3.7$ $68.0\pm 2.2$ $72.0\pm 3.7$ $68.3\pm 3.3$ Assembly -- $67.3\pm 3.7$ $61.7\pm 2.9$ $7.3\pm 0.9$ $35.3\pm 2.5$ $55.7\pm 2.5$ $72.7\pm 3.8$ $82.3\pm 0.5$ $90.3\pm 3.1$ $94.3\pm 0.5$ $5.0\pm 3.6$ $97.0\pm 2.2$ $93.3\pm 1.2$ $92.3\pm 1.7$ $93.0\pm 2.4$triangledown $24.7\pm 0.5$ $47.3\pm 2.9$ $59.7\pm 2.1$ $34.3\pm 2.5$ $39.7\pm 3.8$ $63.3\pm 0.9$ $50.3\pm 2.1$ $82.3\pm 4.7$ $93.0\pm 2.2$ $35.7\pm 29.8$ $77.0\pm 2.9$ $98.3\pm 2.4$ $92.7\pm 2.1$ $93.3\pm 1.2$diamond $73.0\pm 4.1$ $79.3\pm 3.1$ $66.0\pm 4.3$ $53.7\pm 1.9$ $79.3\pm 2.1$ $61.3\pm 2.5$ $63.3\pm 3.1$ $90.3\pm 4.5$ $91.0\pm 2.8$ $71.7\pm 8.7$ $93.7\pm 3.4$ $100.0\pm 0.0$ $87.7\pm 3.3$ $92.7\pm 2.9$ Button Press Wall -- $70.3\pm 4.5$ $76.0\pm 2.9$ $49.3\pm 1.9$ $76.7\pm 4.2$ $62.3\pm 9.0$ $57.7\pm 1.2$ $31.0\pm 0.0$ $57.3\pm 0.9$ $69.3\pm 2.6$ $22.3\pm 17.4$ $52.0\pm 7.9$ $63.3\pm 3.9$ $58.7\pm 4.6$ $51.3\pm 5.4$triangledown $59.3\pm 7.0$ $84.7\pm 2.9$ $63.3\pm 4.9$ $67.7\pm 1.2$ $64.7\pm 1.2$ $74.0\pm 5.4$ $65.0\pm 0.8$ $63.7\pm 7.5$ $64.0\pm 4.9$ $65.7\pm 14.8$ $68.0\pm 2.9$ $60.3\pm 2.6$ $62.0\pm 6.5$ $50.0\pm 1.6$diamond $79.0\pm 4.2$ $77.3\pm 2.4$ $85.0\pm 0.0$ $82.0\pm 2.8$ $69.0\pm 4.3$ $74.3\pm 0.9$ $63.0\pm 3.7$ $77.7\pm 8.7$ $64.7\pm 3.4$ $6.3\pm 2.5$ $83.3\pm 4.8$ $54.7\pm 1.2$ $58.0\pm 5.7$ $67.0\pm 12.7$ Pick Place Wall -- $39.7\pm 7.4$ $29.3\pm 4.5$ $41.0\pm 4.5$ $36.0\pm 5.7$ $59.3\pm 7.3$ $25.7\pm 1.7$ $49.7\pm 7.1$ $84.0\pm 1.6$ $69.3\pm 1.7$ $54.0\pm 15.7$ $61.7\pm 5.4$ $67.0\pm 4.9$ $80.7\pm 0.5$ $76.3\pm 2.4$triangledown $47.0\pm 6.2$ $69.3\pm 5.3$ $52.0\pm 4.2$ $49.0\pm 2.2$ $64.3\pm 3.3$ $58.7\pm 2.1$ $44.3\pm 2.6$ $88.7\pm 2.4$ $89.0\pm 0.8$ $86.3\pm 2.5$ $59.0\pm 9.6$ $95.3\pm 0.5$ $90.0\pm 0.8$ $95.0\pm 2.4$diamond $24.7\pm 2.5$ $59.0\pm 7.3$ $26.7\pm 4.5$ $29.7\pm 4.2$ $40.0\pm 2.2$ $24.3\pm 0.9$ $26.0\pm 4.3$ $79.7\pm 0.9$ $62.0\pm 3.7$ $86.7\pm 2.6$ $49.7\pm 2.6$ $66.0\pm 0.8$ $83.3\pm 5.2$ $73.3\pm 4.6$ Drawer Open -- $90.7\pm 1.2$ $92.0\pm 0.0$ $92.7\pm 1.2$ $87.7\pm 0.9$ $89.3\pm 0.5$ $96.3\pm 1.2$ $95.7\pm 1.7$ $93.3\pm 1.9$ $93.3\pm 1.9$ $95.7\pm 1.7$ $100.0\pm 0.0$ $90.3\pm 1.2$ $95.0\pm 0.8$ $92.0\pm 0.0$triangledown $93.0\pm 1.4$ $93.3\pm 1.9$ $94.0\pm 1.6$ $90.3\pm 1.2$ $89.3\pm 0.5$ $99.7\pm 0.5$ $100.0\pm 0.0$ $95.0\pm 0.8$ $91.7\pm 0.5$ $99.3\pm 0.9$ $100.0\pm 0.0$ $100.0\pm 0.0$ $95.0\pm 0.8$ $92.0\pm 0.0$diamond $100.0\pm 0.0$ $100.0\pm 0.0$ $100.0\pm 0.0$ $99.3\pm 0.9$ $99.3\pm 0.9$ $100.0\pm 0.0$ $100.0\pm 0.0$ $95.7\pm 0.5$ $100.0\pm 0.0$ $95.7\pm 1.7$ $98.0\pm 0.8$ $100.0\pm 0.0$ $100.0\pm 0.0$ $94.7\pm 0.9$ Average -- $52.7$ $51.2$ $40.9$ $48.6$ $48.7$ $52.6$ $48.6$ $65.3$ $67.4$ $50.6$ $60.1$ $67.5$ $66.3$ $65.1$triangledown $46.9$ $57.0$ $52.8$ $48.9$ $49.2$ $55.7$ $52.9$ $65.3$ $66.8$ $54.7$ $57.6$ $71.4$ $69.9$ $67.7$diamond $58.4$ $61.9$ $56.0$ $56.4$ $58.5$ $54.2$ $52.8$ $71.2$ $68.5$ $65.1$ $63.3$ $75.3$ $74.4$ $70.9$